# Part-Time Penalties and Heterogeneous Retirement Decisions

### v1

@@ 3.6 Representative Firm
Nonlinear labor and linear labor for occupations are denoted by $L_{1}$ and $L_{2}$, respectively, while K represents the firm’s capital demand.
@@ 3.7 Stationary Competitive Equilibrium
I assume that the initial asset $a_{0}$ is zero, and I define individual states, $s_{Y}$ and $s_{O}$, as $s_{Y}\equiv(a,x_{-1},j_{-1},\phi,\eta_{1},t,e_{-1},p)\in\textbf{S}^{Y}$ and $s_{O}\equiv(a,t,e_{-1},p)\in\textbf{S}^{O}$, respectively, where the state spaces, $\textbf{S}^{Y}$ and $\textbf{S}^{O}$, are $\textbf{S}^{Y}\equiv\textbf{A}\times\textbf{X}\times\textbf{J}\times\Phi\times\textbf{$\eta$}\times\textbf{T}\times\textbf{E}\times\textbf{P}$ and $\textbf{S}^{O}\equiv\textbf{A}\times\textbf{T}\times\textbf{E}\times\textbf{P}$.
[…]
@@
    Given ($r,w_{1},w_{2}$), policy functions $\{c_{Y}(s_{Y}),a_{Y}^{\prime}(s_{Y}),h_{Y}(s_{Y}))\}_{s_{Y}\in\textbf{S}^{Y}}$, $\{c_{O}(s_{O}),a_{O}^{\prime}(s_{O})\}_{s_{O}\in\textbf{S}^{O}}$ and $\{n(s_{Y}),j(s_{Y})\}_{s_{Y}\in\textbf{S}^{Y}}$ and value functions $\{V^{Y}(s_{Y}),N(s_{Y}),W(s_{Y})\}_{s_{Y}\in\textbf{S}^{Y}}$ and $\{V^{O}(s_{O})\}_{s_{O}\in\textbf{S}^{O}}$ solve the household problem.
@@ 4.3 Source of nonlinearity
1616footnotetext: NL/L:Population ratio is the ratio of the number of Nonlinear workers to that of Linear workers. Also, NL/L:Wage ratio means the ratio of the average efficiency wage of Nonlinear workers to that of Linear workers.
@@ 5 Counterfactual Experiment
First, as shown in Table 4, eliminating the earnings test positively impacts consumption equivalence: the short-term consumption equivalence (CEV)171717 The short-term CEV refers to the consumption equivalent variation for individuals aged 25 immediately following the policy reform, based on the distribution of agents in the baseline model. In contrast, the long-term CEV represents the consumption equivalent variation for individuals of the same age between the two stationary equilibria For example, CEV at age t, denoted as $CEV_{t}$, is defined as follows: Let $\{c_{s}^{0},h_{s}^{0},a_{s+1}^{0}\}_{s=t}^{T}$ and $\{c_{s}^{1},h_{s}^{1},a_{s+1}^{1}\}_{s=t}^{T}$ represent the household’s allocation in an equilibrium of the baseline model and the compared allocation, respectively. $u\Big{(}(1+CEV_{t})c_{t}^{0},h_{t}^{0}\Big{)}+\Big{(}1-S(t)\Big{)}\mu(a_{t+1}^{0})+\sum_{s=t+1}^{T}\beta^{s-t}S(s-1)\bigg{\{}u\Big{(}(1+CEV_{t})c_{s}^{0},h_{s}^{0}\Big{)}+\Big{(}1-S(s)\Big{)}\mu(a_{s+1}^{0})\bigg{\}}$ $=u\Big{(}c_{t}^{1},h_{t}^{1}\Big{)}+\Big{(}1-S(t)\Big{)}\mu(a_{t+1}^{1})+\sum_{s=t+1}^{T}\beta^{s-t}S(s-1)\bigg{\{}u\Big{(}c_{s}^{1},h_{s}^{1}\Big{)}+\Big{(}1-S(s)\Big{)}\mu(a_{s+1}^{1})\bigg{\}}$ and long-term CEV at 25 years old are $0.001061\%$ and $0.01940\%$, respectively. Across all generations, the short-term CEV remains positive but small, as shown in Figure 8. This reform does not affect the extensive margin of older individuals but increases the intensive margin by $2.539\%$. Additionally, it boosts aggregate output, labor supply, capital supply, and tax revenue, though these increases remain under $1\%$.
@@
1717footnotetext: The CEV for the ’Extend Pension Age’ policy is plotted only up to age 64 because individuals aged 65 to 69 lose pension eligibility under this reform, which is embedded within the value functions.
@@ 6 Conclusion
It also demonstrates that while reducing pension benefits decreases labor supply in nonlinear occupations, other policies effectively increase both the intensive and extensive margins of older workers. These heterogeneous effects arise from the interplay of income and substitution effects across different occupational types. When making policy plans, it is important to consider these heterogeneous effects.
For future research, it is essential to investigate additional factors influencing age-friendliness across occupations, such as working conditions, physical strain, and firms’ profit-maximizing strategies. A more comprehensive assessment of these aspects could guide the creation of more effective and targeted policies.
Moreover, examining the labor supply of older females has become increasingly important. While 43.72% of females aged 25–59 are employed part-time, their labor force participation rate remains at approximately 80% within the working-age population. As female workers continue to constitute a growing share of the labor market and approach retirement, understanding their retirement behavior alongside that of males is essential for developing inclusive and comprehensive labor market policies.
@@ A.1. Data Description in Japan
1818footnotetext: ”Leave” represents the proportion of workers who exit the labor force after leaving their occupation. ”Switch across” refers to transitions between different occupation types; for example, a nonlinear worker moving to a linear occupation. ”Switch within” indicates job changes within the same occupation category; for instance, a nonlinear worker moving to another nonlinear job.
@@ A.7. Counterfactual Experiment(Unconventional Policy)
1818footnotetext: When I reduce $\theta_{NL}$ to $\theta_{L}$, since both of $\frac{K}{L}$ and $\frac{EL}{E_{1}}{L_{1}}$ decrease, leading to a reduction in $w_{1}$, which represents the payment per unit of efficiency labor in nonlinear occupations. However, nonlinear workers with high labor disutility increase their productivity and receive a higher hourly wage, while those with lower labor disutility reduce their working hours due to the weakened nonlinearity. As a result, the overall wage level for a nonlinear worker increases.

### v2

@@ 3.6 Representative Firm
Labor inputs for nonlinear and linear occupations are denoted by $L_{1}$ and $L_{2}$, respectively, and are assumed to be complementary. The firm’s capital demand is represented by $K$.
@@ 3.7 Stationary Competitive Equilibrium
The model follows a standard life-cycle framework. The initial asset $a_{0}$ is assumed to be zero, and individual states, $s_{Y}$ and $s_{O}$, are defined as $s_{Y}\equiv(a,x_{-1},j_{-1},\phi,\eta_{1},t,e_{-1},p)\in\textbf{S}^{Y}$ and $s_{O}\equiv(a,t,e_{-1},p)\in\textbf{S}^{O}$, respectively, where the state spaces, $\textbf{S}^{Y}$ and $\textbf{S}^{O}$, are $\textbf{S}^{Y}\equiv\textbf{A}\times\textbf{X}\times\textbf{J}\times\Phi\times\textbf{$\eta$}\times\textbf{T}\times\textbf{E}\times\textbf{P}$ and $\textbf{S}^{O}\equiv\textbf{A}\times\textbf{T}\times\textbf{E}\times\textbf{P}$.
The equilibrium follows a standard definition, where both the capital and two labor markets are clear. In detail, the stationary competitive equilibrium consists of factor prices $(r,w_{1},w_{2})$, allocations to agents,$\{c_{Y}(s_{Y}),a_{Y}^{\prime}(s_{Y}),h_{Y}(s_{Y})\}_{s_{Y}\in\textbf{S}^{Y}}$ and $\{c_{O}(s_{O}),a_{O}^{\prime}(s_{O})\}_{s_{O}\in\textbf{S}^{O}}$, working decision rules of the agents $\{n(s_{Y}),j(s_{Y})\}_{s_{Y}\in\textbf{S}^{Y}}$, value functions, $\{V^{Y}(s_{Y}),N(s_{Y}),W(s_{Y})\}_{s_{Y}\in\textbf{S}^{Y}}$ and $\{V^{O}(s_{O})\}_{s_{O}\in\textbf{S}^{O}}$ allocations to firms $(K^{D},L_{1}^{D},L_{2}^{D})$, and probability measures , $F_{Y}(\cdot)$, on the Borel set $\mathrm{B}(\textbf{S}^{Y})$ such that $F_{Y}(\cdot):\mathcal{B}(\textbf{S}^{Y})\to[0,1]$ and , $F_{O}(\cdot)$, on the Borel set $\mathcal{B}(\textbf{S}^{O})$ such that $F_{O}(\cdot):\mathcal{B}(\textbf{S}^{O})\to[0,1]$.
@@
    Given ($r,w_{1},w_{2}$), policy functions $\{c_{Y}(s_{Y}),a_{Y}^{\prime}(s_{Y}),h_{Y}(s_{Y})\}_{s_{Y}\in\textbf{S}^{Y}}$, $\{c_{O}(s_{O}),a_{O}^{\prime}(s_{O})\}_{s_{O}\in\textbf{S}^{O}}$ and $\{n(s_{Y}),j(s_{Y})\}_{s_{Y}\in\textbf{S}^{Y}}$ and value functions $\{V^{Y}(s_{Y}),N(s_{Y}),W(s_{Y})\}_{s_{Y}\in\textbf{S}^{Y}}$ and $\{V^{O}(s_{O})\}_{s_{O}\in\textbf{S}^{O}}$ solve the household problem.
@@ 4.3 Source of nonlinearity
1414footnotetext: NL/L:Population ratio is the ratio of the number of Nonlinear workers to that of Linear workers. Also, NL/L:Wage ratio means the ratio of the average efficiency wage of Nonlinear workers to that of Linear workers.
@@ 5 Counterfactual Experiment
First, as shown in Table 4, eliminating the earnings test positively impacts consumption equivalence: the short-term consumption equivalence (CEV)151515 The short-term CEV refers to the consumption equivalent variation for individuals aged 25 immediately following the policy reform, based on the distribution of agents in the baseline model. In contrast, the long-term CEV represents the consumption equivalent variation for individuals of the same age between the two stationary equilibria For example, CEV at age t, denoted as $CEV_{t}$, is defined as follows: Let $\{c_{s}^{0},h_{s}^{0},a_{s+1}^{0}\}_{s=t}^{T}$ and $\{c_{s}^{1},h_{s}^{1},a_{s+1}^{1}\}_{s=t}^{T}$ represent the household’s allocation in an equilibrium of the baseline model and the compared allocation, respectively. $u\Big{(}(1+CEV_{t})c_{t}^{0},h_{t}^{0}\Big{)}+\Big{(}1-S(t)\Big{)}\mu(a_{t+1}^{0})+\sum_{s=t+1}^{T}\beta^{s-t}S(s-1)\bigg{\{}u\Big{(}(1+CEV_{t})c_{s}^{0},h_{s}^{0}\Big{)}+\Big{(}1-S(s)\Big{)}\mu(a_{s+1}^{0})\bigg{\}}$ $=u\Big{(}c_{t}^{1},h_{t}^{1}\Big{)}+\Big{(}1-S(t)\Big{)}\mu(a_{t+1}^{1})+\sum_{s=t+1}^{T}\beta^{s-t}S(s-1)\bigg{\{}u\Big{(}c_{s}^{1},h_{s}^{1}\Big{)}+\Big{(}1-S(s)\Big{)}\mu(a_{s+1}^{1})\bigg{\}}$ and long-term CEV at 25 years old are $0.001061\%$ and $0.01940\%$, respectively. Across all generations, the short-term CEV remains positive but small, as shown in Figure 8. This reform does not affect the extensive margin of older individuals but increases the intensive margin by $2.539\%$. Additionally, it boosts aggregate output, labor supply, capital supply, and tax revenue, though these increases remain under $1\%$.
@@
1515footnotetext: The CEV for the ’Extend Pension Age’ policy is plotted only up to age 64 because individuals aged 65 to 69 lose pension eligibility under this reform, which is embedded within the value functions.
@@ 6 Conclusion
This study also demonstrates that while reducing pension benefits decreases labor supply in nonlinear occupations, tax-based reforms—such as increasing tax credits and exempting pension benefits from income taxation—effectively enhance both the intensive and extensive labor supply margins among older workers. These heterogeneous effects stem from the interplay between income and substitution effects across different occupational types. Therefore, policy design should carefully account for such occupational heterogeneity in behavioral responses.
For future research, it is important to examine additional factors that influence age-friendliness across occupations, such as working conditions and health status. It is also necessary to justify the use of linear or nonlinear occupational structures on the production side. Conceptually, workers in nonlinear occupations tend to earn a premium for working full-time because their tasks—such as responding to client calls at any time or attending meetings—are difficult for firms to reallocate or substitute. A more comprehensive assessment of these dimensions could inform the design of more effective and targeted policy interventions.
Moreover, examining the labor supply of older female workers has become increasingly important. In 2023, while 45.56% of women aged 25–59 were employed part-time, their labor force participation rate stood at approximately 81.91%. As women continue to comprise a growing share of the labor market and approach retirement age, understanding their retirement behavior alongside that of men is essential for designing inclusive and comprehensive labor market policies.
@@ A.1. Data Description in Japan
1616footnotetext: ”Leave” represents the proportion of workers who exit the labor force after leaving their occupation. ”Switch across” refers to transitions between different occupation types; for example, a nonlinear worker moving to a linear occupation. ”Switch within” indicates job changes within the same occupation category; for instance, a nonlinear worker moving to another nonlinear job.
@@ A.7. Counterfactual Experiment(Unconventional Policy)
1616footnotetext: When I reduce $\theta_{NL}$ to $\theta_{L}$, since both of $\frac{K}{L}$ and $\frac{EL}{E_{1}}{L_{1}}$ decrease, leading to a reduction in $w_{1}$, which represents the payment per unit of efficiency labor in nonlinear occupations. However, nonlinear workers with high labor disutility increase their productivity and receive a higher hourly wage, while those with lower labor disutility reduce their working hours due to the weakened nonlinearity. As a result, the overall wage level for a nonlinear worker increases.
@@
Table 6: Nonlinear/Linear Occupation List in Japan
No. Occupation (Small Category) Nonlinear/Linear
1 Database SE Nonlinear
2 Pharmaceutical Sales Nonlinear
3 Banking Sales Nonlinear
4 Research and Development(Chemistry) Nonlinear
5 Product Development, Merchandiser Nonlinear
6 Infrastructure Engineer Nonlinear
7 Customer Engineer Nonlinear
8 Machinery Sales Nonlinear
9 System, IT Consultant Nonlinear
10 Public Health Nurse, Midwife Nonlinear
11 Human Resources Nonlinear
12 Web Designer Nonlinear
13 Research and Development(Biotechnology) Nonlinear
14 Technical Development(Construction, Civil Engineering, Plant, Equipment) Nonlinear
15 Doctor, Dentist, Veterinarian Nonlinear
16 Marketing Nonlinear
17 Business Planning Nonlinear
18 Store Development, Other Planning, Promotion Office Professional Nonlinear
19 Pharmacist Nonlinear
20 Purchasing, Materials Nonlinear
21 Communication, Network Engineer Nonlinear
22 Railway Operator, Telephone Operator, Mail Delivery, etc. Nonlinear
23 Programmer Nonlinear
24 Civil Engineering Design Nonlinear
25 Manager(Professional) Nonlinear
26 Other Financial Specialist Occupation Nonlinear
27 Manager(Service) Nonlinear
28 Planning, Sales Promotion Nonlinear
29 Support Engineer(Software) Nonlinear
30 Legal Affairs Nonlinear
31 Other Internet-related Technical Occupation Nonlinear
32 Other SE Nonlinear
33 Teacher, Lecturer, Instructor, Interpreter, etc. Nonlinear
34 Manager(Administrative) Nonlinear
35 Nurse(including Assistant Nurse) Nonlinear
36 Radiologic Technologist, Clinical Laboratory Technician, Dental Technician, Physical Therapist, etc. Nonlinear
37 Other Office Worker Nonlinear
38 Web Application Development Nonlinear
39 General Affairs Nonlinear
40 Food Sales Nonlinear
41 Author, Journalist, Editor, Proofreader, etc. Nonlinear
42 Public Relation Nonlinear
43 Business Nonlinear
44 Inventory Management Nonlinear
45 Electrical Circuit Design Nonlinear
46 Research and Development(Food) Nonlinear
47 Product Management Nonlinear
48 Other Research and Development Nonlinear
49 Welfare Counseling Specialist, Childcare Worker, Caregiver, etc. Nonlinear
50 Other Advertising, Publishing, Media-Related Professional Position Nonlinear
51 Receptionist Nonlinear
52 Product Planning Nonlinear
53 Architectural Design Nonlinear
54 Other Technical Occupation Nonlinear
55 Research and Development(Machinery) Nonlinear
56 Administrative Management Nonlinear
57 Metal, Machinery, Electrical, Automobile Manufacturing, Production, Repair Worker Nonlinear
58 Real Estate Sales Nonlinear
59 Medical Administration Nonlinear
60 Other Building, Civil Engineering, Surveying Technicians Nonlinear
61 Self-Defense Officer, Police Officer, Security Guard, etc. Nonlinear
62 Web Producer, Director, Planner Nonlinear
63 Telecommunications Technician Nonlinear
64 Sales Administration Nonlinear
65 Other General Office Professional Nonlinear
66 Development(Software-related Occupation) Linear
67 Unclassified Occupation Linear
68 Certified Public Accountant, Tax Accountant, etc. Linear
69 Keypuncher, Computer Operator, etc. Linear
70 Other Unclassified Service Professional Linear
71 Other Manufacturing Worker Linear
72 Cleaning, Delivery, Warehouse Work, etc. Linear
73 Finance, Accounting Linear
74 Manager(Business) Linear
75 Manager(Sales) Linear
76 Service Staff(Gas Station) Linear
77 Labor Affairs Linear
78 Other Electrical, Electronic, Mechanical Design-Related Professions Linear
79 Building, Parking Lot, Condominium, Boiler Management Linear
80 Nutritionist, Masseur, Counselor, etc. Linear
81 Other Customer Service and Serving Occupation Linear
82 Lawyer, Patent Agent, Judicial Scrivener, etc. Linear
83 Secretary Linear
84 Other Engineer Linear
85 Housekeeper, Home Helper, etc. Linear
86 Telephone Operator Linear
87 Telecommunications Sales Linear
88 DTP Operator Linear
89 Other Sales Linear
90 Store Clerk, Fashion Advisor, Cashier Linear
91 Food, Daily Necessities Manufacturing, Production Worker Linear
92 Driver(Van, Wagon) Linear
93 Other Cooking Professionals, Bartender Linear
94 Other Printing-related Specialist Occupation Linear
95 Store Manager Linear
96 Other Life and Hygiene Service Professional Linear
97 Other Construction, Civil Engineering, Mining Worker Linear
98 Manager(Technical) Linear
99 Management and Accounting Consultant, etc. Linear
100 Facility Construction Site Management, Site Supervisor, Construction Management Linear
101 Systems Sales Linear
102 Arrangement Operation Linear
103 Procurement Linear
104 Fashion-related Occupation Linear
105 Driver(Bus) Linear
106 Manager(Other) Linear
107 Mechanical Design Linear
108 Driver(Truck) Linear
109 Research and Development(Electrical, Electronic) Linear
110 Accommodation Services Linear
111 Machinery Maintenance Linear
112 Waiter, Waitress Linear
113 Farming, Landscaping, Livestock Worker, Forestry, Fisheries Worker Linear
114 Insurance Sales Linear
115 Construction Worker(Construction Worker) Linear
116 Real Estate Mediator, Salesperson, Insurance Agent, etc. Linear
117 Photographer Linear
118 Barber, Beautician Linear
119 Esthetician Linear
120 Driver(Taxi, Limousine) Linear
121 Western Cuisine Chef Linear
122 Construction Worker(Facility Construction Worker) Linear
123 Civil Engineering Site Management, Site Supervisor, Construction Supervision Linear
124 Other Design Linear
125 Electrical, Electronic Equipment Sales Linear
126 Automobile, Motorcycle Mechanic Linear
127 Hall Staff(Pachinko, Amusement Arcade) Linear
128 Trade Administration Linear
129 Japanese Cuisine Chef, Sushi Chef Linear
130 Construction Site Management, Site Supervisor, Construction Supervision Linear
131 Supervisor Linear
132 Character, CG Designer Linear
133 Construction Worker(Civil Engineer) Linear
134 Control SE Linear
135 Control Design Linear